# Capsule Network Projectors are Equivariant and Invariant Learners

### v2

@@ Abstract
Learning invariant representations has been the longstanding approach to self-supervised learning. However, recently progress has been made in preserving equivariant properties in representations, yet do so with highly prescribed architectures. In this work, we propose an invariant-equivariant self-supervised architecture that employs Capsule Networks (CapsNets) which have been shown to capture equivariance with respect to novel viewpoints. We demonstrate that the use of CapsNets in equivariant self-supervised architectures achieves improved downstream performance on equivariant tasks with higher efficiency and fewer network parameters. To accommodate the architectural changes of CapsNets, we introduce a new objective function based on entropy minimisation. This approach which we name CapsIE (Capsule Invariant Equivariant Network) achieves state-of-the-art performance across invariant and equivariant tasks on the 3DIEBench dataset compared to prior equivariant SSL methods, while outperforming supervised baselines. Our results demonstrate the ability of CapsNets to learn complex and generalised representations for large-scale, multi-task datasets compared to previous CapsNet benchmarks. Code is available at https://github.com/AberdeenML/CapsIE.
@@ 1 Introduction
Equivariance and invariance have become increasingly important properties and objectives of deep learning in recent times, with precedence being largely placed on the latter. The task of invariance, i.e. being able to classify a specific object no matter the camera perspective or augmentation applied, has driven progress in modern self-supervised learning approaches, specifically those which follow a joint embedding architecture [1, 3, 7]. Equivariance on the other hand is the task of capturing embeddings which equally reflect the translations applied to the input space in the latent space. Equivariance thus has become an important property to capture to enable the learning of high-quality representations in the real world where transformations such as viewpoint are essential.
@@
Self-supervised learning predominantly owes its success to such invariant objectives, where all recent progress, whether that is by contrastive [7], information-maximisation [3, 33], or clustering based methods [5, 1] rely on ensuring invariance in their representations under augmentation. This setting typically ensures performance in classification based tasks, but when employing the representations in alternative tasks, preservation of information is essential to improve generalisation. To maintain properties of the transformation one can predict the augmentations applied [8, 21], yet this is typically not considered truly equivariant given the mapping of transformations is not represented in the latent space. Methods that employ such a prediction methodology are typically considered equivariant as the transformation in the input space is directly preserved in the latent space. Here prediction networks are employed to reconstruct the view prior to transformation [30], learn symmetric representations [26], or predict the latent representation of the transformed view from the representation of the original view given the transformation parameters [14].
@@
The above methods, although promising, enforce equivariance via objective functions on vector representations, yet these methods fail to employ architectural approaches that have shown to be capable of better capturing these properties. Capsule Networks (CapsNets), which utilise a process called routing [29, 17, 12, 15, 9, 22], are one such architecture, showing signs of desirable properties that other state of the art (SOTA) architectures such as Vision Transformers (ViTs) and CNNs do not. Specifically, CapsNets have shown a natural ability to have strong viewpoint equivariance and viewpoint invariance properties – they achieve this through their ability to capture equivariance with respect to viewpoints in neural activities, and invariance in the weights. In addition, viewpoint changes have nonlinear effects on pixels but linear effects on object relationships [9, 17]. Ideally, these properties could lead to the development of more sample-efficient models that can exploit robust representations to better generalise to unseen cases and new samples.
@@
However, a common argument is that CapsNets have only shown these properties on toy examples such as the smallNORB dataset [20], which many would consider irrelevant for modern architectures. Despite this, small CapsNets outperform much larger CNN and ViT counterparts [12]. In this work, we propose a novel CapsNet formulation and corresponding objective function, achieving SOTA on several experiments and ablations studies on the 3DIEBench dataset [14] which has been created to specifically benchmark equivariant and invariant properties of deep learning models. We prove that CapsNets retain their desirable properties on this dataset which is considerably more difficult than what has been previously achieved with CapsNets, while also establishing new SOTA on these tasks.
@@
Figure 1: Left: Schematic overview of the proposed CapsIE architecture. Representations are fed into a CapsNet projector, and the output embeddings $Z_{\text{act}}$ and $Z_{\text{pose}}$ correspond to invariant and equivariant embeddings respectively. Right: Generalised view of a Capsule projection head. A CNN feature map is transformed via the primary capsules into initial poses $u_{i}$, represented by cylinders, and activations $a_{i}$, represented by circles. These poses are transformed into votes, which represent a lower-level capsules prediction for each of the higher-level capsules. The routing process then defines how well these votes match the concept represented by the upper-level capsule, creating the coupling coefficients. Finally, these coupling coefficients are used to create $u_{j}$ and $a_{j}$, the output of the capsule projector head.
@@
  • We demonstrate that CapsNet projectors implicitly learn pose understanding in a self-supervised setting.
  • We show state of the art performance on 3DIEBench classification for equivariance and invariance benchmark tasks from our CapsNet based architecture.
@@
2 Problem Statement
Typically, self-supervised learning maximises the similarity between embeddings of two augmented views of an image such that they are invariant to augmentations, and instead capture semantically meaningful information of the original image. Views $x$ and $x^{\prime}$ are each transformed from an image $d\in\mathbb{R}^{c\times h\times w}$ sampled from dataset $\mathcal{D}$ by image augmentations $\tau,\tau^{\prime}\sim T$ sampled from a set of augmentations $T$. Embeddings are obtained by feeding each view through an encoder $f_{\theta}$, where the output representations $y,y^{\prime}$ are fed through a projection head $h_{\phi}$ to produce embeddings $z,z^{\prime}$ whose similarity is maximised. However, it is detrimental in many settings that $f$ is invariant to all transformations, instead in this work we are focused on ensuring that $f$ is equivariant to viewpoint transformations. To train for and evaluate such properties we base our study on the challenging 3DIEBench dataset and corresponding problem definition presented in [14].
First to define equivariance we begin by defining a Group consisting of a set $G$ and a binary operation $\cdot$ on $G$, $\cdot:G\times G\rightarrow G$ such that $\cdot$ are associative; there is an identity $e$ which satisfies $e\cdot a=a=a\cdot e,\forall a\in G$; and for each $a\in G$ there exists an inverse $a^{-1}$ such that $a\cdot a^{-1}=e=a^{-1}\cdot a$. Group actions are concerned with how groups manipulate sets, where the left group action can be defined as a function $\alpha$ of group $G$ and set $S$, $\alpha:G\times S\rightarrow S$ such that $\alpha(e,s)=s,\forall s\in S$, and $\alpha(g,\alpha(h,s))=\alpha(gh,s),\forall s\in S,\text{and }\forall g,h\in G$. In our setting, we are concerned with group representations which are linear group actions acting on vector space $V$, which we define as $\rho:G\rightarrow GL(V)$ where $GL(V)$ is the general linear group on $V$. Here $\rho(g)$ describes the transformation applied to both the input data $x$ and latent $f(x)$ given parameters $g$ [26, 14].
@@ 2 Problem Statement
As defined in [14] the goal (visually depicted in Figure 2) is to learn $f$ and $\rho_{Y}$ to construct representations that are equivariant to viewpoint transformations when $\rho_{X}$ is not known, but the group elements $g$ that parameterise the transformations are known. Further details of these transformations and the benchmark 3DIEBench dataset are given in Section A.
@@ 3.1 Architecture
Our method which we name CapsIE (Capsule Network Invariant Equivariant) follows the general joint embedding architecture previously described in Section 2 and extends those proposed by VICReg [3] and SIE [14]. Like previous methods, we employ a ResNet-18 [16] encoder as the core feature extractor $f_{\theta}$ of our network. Yet unlike SIE [14], we do not split the representations, and therefore do not require the use of separate invariant and equivariant projection heads. Instead, we employ a single CapsNet (described in Section 3.2) which takes as input the full representation of the encoder in place of the multi-layer perceptron (MLP) in the projection head $h_{\phi}$. Given the architectural design of CapsNets, our projection head outputs both an activation scalar, representing how active the capsule is, and a $4\times 4$ pose for each capsule.
To align with the above problem statement and Equation 1, we aim to simultaneously learn invariant and equivariant representations by optimising our network $f$ with respect to the output activations and poses. In this case, we consider the activation vector to capture existence of semantic concepts/objects of the input, thus, the invariant information is preserved by the transformation. The pose on the other hand is designed to encode positional information related to each corresponding capsule [27] (i.e. semantic concept), therefore it contains equivariant information that that was changed by the transformation. Akin to the SIE method, we therefore, consider two embedding vectors for each image view, $z_{\text{act}}$ and $z_{\text{pose}}$, which correspond to the capsule activation vector and pose, and invariant and equivariant components respectively.
To enforce our network to learn equivariant properties we utilise a prediction network $p_{\psi,g}$ which takes as input the transformation $g$ and $z_{\text{pose}}$ to predict $z^{\prime}_{\text{pose}}$, and hence learn $\rho_{Y}(g)$. In our setting, $g\in\mathbb{R}^{3}$ corresponding to the quaternions of the rotation applied. In this work, we employ the hypernetwork approach taken by [14] which uses an linear projector that takes as input the transformation parameters $g$ to parameterise an MLP predictor. Such a network avoids the case where the transformation parameters $g$ are ignored and the predictor provides invariant solutions. We present more details of the predictor network in the appendix, and visually depict the full CapsIE architecture in Figure 1.
@@ 3.2 Capsule Network Projector
Figure 3: Simplified visual representation of CapsNet outputs. Activation vector outputs probability of each capsule being activated. pose corresponds to the object pose for each corresponding capsule.
@@
CapsNets are designed to handle spatial hierarchies and recognise objects regardless of their orientation or location, achieving equivariance through their structure [28]. A capsule is a group of neurons – vector-based representations – representing instantiation parameters such as position, orientation, and size. Before any routing process begins, lower-level capsule poses $u_{i}$ are transformed to $n$ upper-level capsule poses $u_{j|i}$ which align with concepts represented by higher-level capsules, preserving spatial relationships and hierarchical information. It is then determined through the routing process how well these transformed poses correspond with the concept represented by the upper-level capsule. CapsNets, unlike convolution, excel in achieving viewpoint invariance and viewpoint equivariance as they can capture equivariance with respect to viewpoints in neural activities, and invariance in the network’s weights [27]. Consequently, capsule routing aims to detect objects by looking for agreement between their parts, thereby performing equivariant inference.
@@
We use the Self Routing CapsNet (SRCaps) [15] based on the efficiency of its non-iterative routing algorithm. We consider the trade-off of a small amount of classification accuracy to be acceptable when comparing the performance of SRCaps to other capsule architectures which require significantly more resources to train. Based on the size of the 3DIEBench dataset, these other routing algorithms would be unsuitable.
@@ 3.3 Objective Functions
To train our aforementioned architecture we first introduce an invariant objective as the cross entropy between activation probability vectors, $H(Z_{\text{act}},Z^{\prime}_{\text{act}})$ where $Z$ refers to the matrix embeddings over a batch. The aims is to enforce embedding probability pairs originating from the same image to be matched. To avoid trivial solutions and collapse to a single capsule, we employ the mean entropy maximisation regularisation [2, 1] on the same activation probability vectors to encourage the model to utilise the full set of capsules over a batch. This regularisation maximises the entropy of the mean probabilities $H(\bar{Z}_{\text{act}})$ and $H(\bar{Z}^{\prime}_{\text{act}})$, where $\bar{Z}_{\text{act}}=\frac{1}{B}\sum^{B}_{i=1}Z_{\text{act}}$ and $B$ is the batch size.
@@
As previously stated in Section 2, our goal is to learn the predictor $p_{\psi,g}$ to model $\rho_{Y}(g)$ as to enforce equivariant representations. This is achieved by maximising the cosine similarity between the output vector of the predictor $p_{\psi,g}(Z_{\text{pose}})$ given translation parameters $g$ and equivariant representation $Z_{\text{pose}}$, and the augmented view’s equivariant representation vector $Z^{\prime}_{\text{pose}}$. To avoid collapse and improve training stability we also regularise the output of $p_{\psi,g}(Z_{\text{pose}})$ by ensuring the variance of the predicted equivariance representation is 1 to avoid collapse. Whereas, SIE [14] finds this to be an optional but recommended component, we found in practice, without such regularisation the predictor would consistently collapse to trivial solutions.
@@
The final objective function is given by the weighted sum of the individual objectives

### v3

@@ Abstract
Learning invariant representations has been the longstanding approach to self-supervised learning. However, recently progress has been made in preserving equivariant properties in representations, yet do so with highly prescribed architectures. In this work, we propose an invariant-equivariant self-supervised architecture that employs Capsule Networks (CapsNets) which have been shown to capture equivariance with respect to novel viewpoints. We demonstrate that the use of CapsNets in equivariant self-supervised architectures achieves improved downstream performance on equivariant tasks with higher efficiency and fewer network parameters. To accommodate the architectural changes of CapsNets, we introduce a new objective function based on entropy minimisation. This approach which we name CapsIE (Capsule Invariant Equivariant Network) achieves state-of-the-art performance on the equivariant rotation tasks on the 3DIEBench dataset compared to prior equivariant SSL methods, while performing competitively against supervised counterparts. Our results demonstrate the ability of CapsNets to learn complex and generalised representations for large-scale, multi-task datasets compared to previous CapsNet benchmarks. Code is available at https://github.com/AberdeenML/CapsIE. Note, this version presents corrected results from the prior version, where a bug in the code had been identified, leading to some minor discrepancies.
@@ 1 Introduction
Equivariance and invariance have become increasingly important properties and objectives of deep learning in recent times, with precedence being largely placed on the latter. The task of invariance, i.e. being able to classify a specific object no matter the camera perspective or augmentation applied, has driven progress in modern self-supervised learning approaches, specifically those which follow a joint embedding architecture (Assran et al., 2022; Bardes et al., 2021; Chen et al., 2020). Equivariance on the other hand is the task of capturing embeddings which equally reflect the translations applied to the input space in the latent space. Equivariance thus has become an important property to capture to enable the learning of high-quality representations in the real world where transformations such as viewpoint are essential.
@@
Self-supervised learning owes its success to invariant objectives, where all recent progress, whether that is by contrastive (Chen et al., 2020), information-maximisation (Bardes et al., 2021; Zbontar et al., 2021), or clustering based methods (Caron et al., 2021; Assran et al., 2022) rely on ensuring invariance in their representations under augmentation. This setting ensures performance in classification based tasks, but when employing the representations in alternative tasks, preservation of information is essential to improve generalisation. To maintain properties of the transformation one can predict the augmentations applied (Dangovski et al., 2022; Lee et al., 2021), yet this is typically not considered truly equivariant given the mapping of transformations is not represented in the latent space. Methods that employ such a prediction methodology are typically considered equivariant as the transformation in the input space is preserved in the latent space. Here prediction networks are employed to reconstruct the view prior to transformation (Winter et al., 2022), learn symmetric representations (Park et al., 2022), or predict the latent representation of the transformed view from the representation of the original view given the transformation parameters (Garrido et al., 2023).
@@
The above methods, although promising, enforce equivariance via objective functions on vector representations, yet these methods fail to employ architectural approaches that have shown to be capable of better capturing these properties. Capsule Networks (CapsNets), which utilise a process called routing (Sabour et al., 2017; Hinton et al., 2018; Everett et al., 2023; Hahn et al., 2019; De Sousa Ribeiro et al., 2020; Liu et al., 2024), are one such architecture, showing signs of desirable properties that other state of the art (SOTA) architectures such as Vision Transformers (ViTs) and CNNs do not. Specifically, CapsNets have shown a natural ability to have strong viewpoint equivariance and viewpoint invariance properties – they achieve this through their ability to capture equivariance with respect to viewpoints in neural activities, and invariance in the weights. In addition, viewpoint changes have nonlinear effects on pixels but linear effects on object relationships (De Sousa Ribeiro et al., 2020; Hinton et al., 2018). Ideally, these properties could lead to the development of more sample-efficient models that can exploit robust representations to better generalise to unseen cases and new samples.
@@
However, a common argument is that CapsNets have only shown these properties on toy examples such as the smallNORB dataset (LeCun et al., 2004), which many would consider irrelevant for modern architectures. Despite this, small CapsNets outperform much larger CNN and ViT counterparts (Everett et al., 2023). In this work, we propose a novel CapsNet formulation and corresponding objective function, achieving SOTA on equivariant viewpoint rotation tasks on the 3DIEBench dataset (Garrido et al., 2023) which has been created to specifically benchmark equivariant and invariant properties of deep learning models. We prove that CapsNets retain their desirable properties on this dataset which is considerably more difficult than what has been previously achieved with CapsNets, while also establishing new SOTA for this dataset.
@@
Figure 1: Left: Schematic overview of the proposed CapsIE architecture. Representations are fed into a CapsNet projector, and the output embeddings $Z_{\text{act}}$ and $Z_{\text{pose}}$ correspond to invariant and equivariant embeddings respectively. Right: Generalised view of a Capsule projection head. CNN feature maps are transformed via the primary capsules into poses $u_{i}$, represented by cylinders, and activations $a_{i}$, represented by circles. Poses are transformed to votes, which represent a lower-level capsules prediction for each of the higher-level capsules. The routing process then defines how well these votes match the concept represented by the upper-level capsule, creating the coupling coefficients. Coupling coefficients inform $u_{j}$ and $a_{j}$, the output of the capsule projector head.
@@
  • We show state of the art performance on 3DIEBench classification for equivariance benchmark tasks from our CapsNet-based architecture.
2 Problem Statement
Typically, self-supervised learning maximises the similarity between embeddings of two augmented views of an image such that they are invariant to augmentations, and instead capture semantically meaningful information of the original image. Views $x$ and $x^{\prime}$ are each transformed from an image $d\in\mathbb{R}^{c\times h\times w}$ sampled from dataset $\mathcal{D}$ by image augmentations $\tau,\tau^{\prime}\sim T$ sampled from a set of augmentations $T$. Embeddings are obtained by feeding each view through an encoder $f_{\theta}$, where the output representations $y,y^{\prime}$ are fed through a projection head $h_{\phi}$ to produce embeddings $z,z^{\prime}$ whose similarity is maximised. However, it is detrimental in many settings that $f$ is invariant to all transformations, instead in this work we are focused on ensuring that $f$ is equivariant to viewpoint transformations. To train for and evaluate such properties we base our study on the challenging 3DIEBench dataset and corresponding problem definition presented in (Garrido et al., 2023).
@@ 2 Problem Statement
First, to define equivariance we begin by defining a Group consisting of a set $G$ and a binary operation $\cdot$ on $G$, $\cdot:G\times G\rightarrow G$ such that $\cdot$ are associative; there is an identity $e$ which satisfies $e\cdot a=a=a\cdot e,\forall a\in G$; and for each $a\in G$ there exists an inverse $a^{-1}$ such that $a\cdot a^{-1}=e=a^{-1}\cdot a$. Group actions are concerned with how groups manipulate sets, where the left group action can be defined as a function $\alpha$ of group $G$ and set $S$, $\alpha:G\times S\rightarrow S$ such that $\alpha(e,s)=s,\forall s\in S$, and $\alpha(g,\alpha(h,s))=\alpha(gh,s),\forall s\in S,\text{and }\forall g,h\in G$. In our setting, we are concerned with group representations which are linear group actions acting on vector space $V$, which we define as $\rho:G\rightarrow GL(V)$ where $GL(V)$ is the general linear group on $V$. Here $\rho(g)$ describes the transformation applied to both the input data $x$ and latent $f(x)$ given parameters $g$ (Park et al., 2022; Garrido et al., 2023). Transformations comprise colour scaling and shifting, and rotations around a fixed point, see Appendix A for further details.
@@
As defined in (Garrido et al., 2023) the goal (visually depicted in Figure 2) is to learn $f$ and $\rho_{Y}$ to construct representations that are equivariant to viewpoint transformations when $\rho_{X}$ is not known, but the group elements $g$ that parameterise the transformations are known. Further details of these transformations and the benchmark 3DIEBench dataset are given in Section A.
@@ 3.1 Architecture
Our method which we name CapsIE (Capsule Network Invariant Equivariant) follows the general joint embedding architecture previously described in Section 2 and extends those proposed by VICReg (Bardes et al., 2021) and SIE (Garrido et al., 2023). Like previous methods, we employ a ResNet-18 (He et al., 2016) encoder as the core feature extractor $f_{\theta}$ of our network. Yet unlike SIE (Garrido et al., 2023), we do not split the representations, and therefore do not require the use of separate invariant and equivariant projection heads. Instead, we employ a single CapsNet (described in Section 3.2) which takes as input the full representation of the encoder in place of the multi-layer perceptron (MLP) in the projection head $h_{\phi}$. Given the architectural design of CapsNets, our projection head outputs both an activation scalar, representing how active the capsule is, and a $4\times 4$ pose for each capsule.
To align with the above problem statement and Equation 1, we aim to simultaneously learn invariant and equivariant representations by optimising our network $f$ with respect to the output activations and poses. In this case, we consider the vector of activations to capture existence of semantic concepts/objects of the input, thus, the invariant information is preserved by the transformation. The pose on the other hand is designed to encode positional information related to each corresponding capsule (Ribeiro et al., 2022) (i.e. semantic concept), therefore it contains equivariant information that that was changed by the transformation. Akin to the SIE method, we therefore, consider two embedding vectors for each image view, $z_{\text{act}}$ and $z_{\text{pose}}$, which correspond to the capsule activation and pose, and invariant and equivariant components respectively.
To enforce our network to learn equivariant properties we utilise a prediction network $p_{\psi,g}$ which takes as input the transformation $g$ and $z_{\text{pose}}$ to predict $z^{\prime}_{\text{pose}}$, and hence learn $\rho_{Y}(g)$. In our setting, $g\in\mathbb{R}^{3}$ corresponding to the quaternions of the rotation applied. In this work, we employ the hypernetwork approach taken by (Garrido et al., 2023) which uses an linear projector that takes as input the transformation parameters $g$ to parameterise an MLP predictor. Such a network avoids the case where the transformation parameters $g$ are ignored and the predictor provides invariant solutions. We present more details of the predictor network in the appendix, and visually depict the full CapsIE architecture in Figure 1.
@@ 3.2 Capsule Network Projector
Figure 3: Simplified visual representation of CapsNet outputs. Vector of activations outputs the probability of each capsule being activated, whereas the pose matrix corresponds to the object pose in relation to the frame.
@@
CapsNets are designed to handle spatial hierarchies and recognise objects regardless of their orientation or location, achieving equivariance through their structure (Ribeiro et al., 2020). A capsule is a group of neurons – vector-based representations – representing instantiation parameters such as position, orientation, and size. Before any routing process begins, lower-level capsule poses $u_{i}$ are transformed to $n$ upper-level capsule poses $u_{j|i}$ which align with concepts represented by higher-level capsules, preserving spatial relationships and hierarchical information. It is then determined through the routing process how well these transformed poses correspond with the concept represented by the upper-level capsule. CapsNets, unlike convolution, excel in achieving viewpoint invariance and viewpoint equivariance as they can capture equivariance with respect to viewpoints in neural activities, and invariance in the network’s weights (Ribeiro et al., 2022). Consequently, capsule routing aims to detect objects by looking for agreement between their parts, thereby performing equivariant inference.
@@
We use the Self Routing CapsNet (SRCaps) (Hahn et al., 2019) based on the efficiency of its non-iterative routing algorithm. We consider the trade-off of a small amount of classification accuracy to be acceptable when comparing the performance of SRCaps to other capsule architectures which require significantly more resources to train. Based on the size of the 3DIEBench dataset, these other routing algorithms would be unsuitable.
@@ 3.3 Objective Functions
To train our aforementioned architecture we first introduce an invariant objective as the cross entropy between activation probability vectors, $H(Z_{\text{act}},Z^{\prime}_{\text{act}})$ where $Z$ refers to the matrix embeddings over a batch. The aim is to enforce embedding probability pairs originating from the same image to be matched. To avoid trivial solutions and collapse to a single capsule, we employ the mean entropy maximisation regularisation (Assran et al., 2021, 2022) on the same activation probability vectors to encourage the model to utilise the full set of capsules over a batch. This regularisation maximises the entropy of the mean probabilities $H(\bar{Z}_{\text{act}})$ and $H(\bar{Z}^{\prime}_{\text{act}})$, where $\bar{Z}_{\text{act}}=\frac{1}{B}\sum^{B}_{i=1}Z_{\text{act}}$ and $B$ is the batch size.
@@
As previously stated in Section 2, our goal is to learn the predictor $p_{\psi,g}$ to model $\rho_{Y}(g)$ as to enforce equivariant representations. This is achieved by maximising the cosine similarity between the output vector of the predictor $p_{\psi,g}(Z_{\text{pose}})$ given translation parameters $g$ and equivariant representation $Z_{\text{pose}}$, and the augmented view’s equivariant representation vector $Z^{\prime}_{\text{pose}}$. To avoid collapse and improve training stability we also regularise the output of $p_{\psi,g}(Z_{\text{pose}})$ by ensuring the variance of the predicted equivariance representation is 1 to avoid collapse. Whereas, SIE (Garrido et al., 2023) finds this to be an optional but recommended component, we found in practice, without such regularisation the predictor would consistently collapse to trivial solutions.
@@
The final objective function is given by the weighted sum of the individual objectives: